OpenStudy (anonymous):

2 over 5 equal x plus 1 over 5

OpenStudy (one098):

\[\frac{ 2 }{ 5 }=\frac{ x+1 }{ 5 }\] Separate the x+1/5 into two fractions. \[\frac{ 2 }{ 5 }=\frac{ x }{ 5 }+\frac{ 1 }{ 5 }\] Now switch the 2/5 to the other side. \[\frac{ x }{ 5 }+\frac{ 1 }{ 5 }=\frac{ 2 }{ 5 }\] Now do the opposite. Subtract 2 - 1 and that's your answer.

OpenStudy (anonymous):

thanks!

OpenStudy (one098):

You're welcome.

OpenStudy (anonymous):